# SelectIT: Selective Instruction Tuning for Large Language Models via Uncertainty-Aware Self-Reflection

### v1

@@ 4.2 Main Results
We focus on the discussion of LLaMA-2-13B because both 7B and 13B models exhibit similar trends in Table 4. System (10) shows the vanilla IT on LLMs with the original Alpaca. By using the data selection strategies, the ability of LLMs has a moderate enhancement in Systems (12) to (14). Additionally, we can use $S^{base}$ as the input for Equations 4 and 5 to construct individual methods of Sentence-R and Model-R. Systems (15) to (17) illustrate that applying each submodule of SelectIT incrementally enhances LLMs’ performance, rivaling contemporary advanced methods.
Most remarkably, SelectIT markedly boosts LLaMA-2’s performance compared to vanilla IT in the System (18). This enhancement is particularly evident in the computational and reasoning tasks on the BBH and GSM benchmarks, compared to other IT data selection strategies. This may be attributed to the characteristics of selected data by SelectIT, and we will analyze this phenomenon in a later section. These gains in reasoning ability also positively impact the coding proficiency of LLMs. The improvement of LLMs on the TydiQA dataset is also obvious enough, which shows that SelectIT can effectively eliminate similar samples and retain sufficient diversity in multilingual aspects.
@@ 5 Analysis
We aim to answer the research questions through the following experiments:
  • How to select high-quality data in SelectIT?
  • Is SelectIT adaptable to various models and domains?
  • What are the advantages of Selective Alpaca?
@@ 5.1 Abalation Study of SelectIT
Figure 3: Comparison of reasoning ability of LLMs with different proportions of Alpaca.
While SelectIT already excels at assessing and ranking samples effectively, selecting an appropriate number of samples in a redundant dataset remains a crucial aspect of our method. We divide the Alpaca dataset into multiple subsets ranging from 10% to 100% based on SelectIT’s evaluation and evaluate the reasoning ability of LLMs which is the SelectIT mainly enhances. As illustrated in Figure 3, compared to the full dataset, we observe that LLMs achieve optimal performance using only 20% of the data. Although the LLaMA-2-7B datasets of 20% and 40% displayed similar capabilities, the former required less fine-tuning time and computational resources. Hence, we opt for the 20% for implementing the SelectIT on the Alpaca dataset.
@@
BBH GSM BBH GSM
3 40.1 17.5 49.6 34.0 35.3
5 40.8 17.5 50.1 34.5 35.7
7 41.2 18.0 49.5 34.0 35.7
9 40.5 17.0 50.3 34.0 35.5
Table 2: Effect of different $K$ on LLMs’ performance
ID Individual Method Unique (%) Overall (%)
@@
Table 3: The relationship between the SelectIT and the individual selection strategy. Sentence-R plays the most significant impact on the final rating of the IT data.
Effect of $K$
$K$ is a critical parameter for our method, impacting not only the range of scores assigned by the LLMs but also the number of rating prompts. We set $K={3,5,7,9}$ and apply the SelectIT algorithm for sample selection within the Alpaca to get different subset datasets. The results in Table 3 indicate that variations in the value of $K$ have a minimal impact on the overall performance of the LLMs. This is attributed to our multi-granularity self-reflection mechanism, which effectively enhances the accuracy and stability of sample selection. Although the model achieves competitive performance at $K$ values of 5 and 7, to minimize resource consumption, we set $K=5$ as the default value in SelectIT.
Datasets Datasets MMLU BBH GSM Tydiqa CodeX AE Overall
AVG $\Delta$ (normal-$\uparrow$)
LLaMA-2-7B Alpaca-GPT4 46.2 39.2 15.0 43.3 27.8 33.7 34.2 -
Selective Alpaca 47.7 40.8 17.5 47.8 29.8 35.4 36.5 +2.3
LLaMA-2-13B Alpaca-GPT4 55.7 47.3 31.0 49.1 41.8 46.5 45.2 -
Selective Alpaca 55.5 50.1 34.5 54.4 43.2 48.2 47.7 +2.5
@@
Table 4: The overall results on IT of different LLMs.
@@
Table 5: The overall results on machine translation LLMs.
Effect of Different Reflection Strategy
We analyze the relationship between individual selection strategies and SelectIT, from the following two aspects. We first account for the number of high-quality data that only can be selected by a specific selection strategy, referred to as unique selection. Secondly, we calculate the percentage of the samples selected by individual selection strategies in Selective Alpaca, called overall selection. As shown in Table 3, Sentence-R plays the most important role in the final SelectIT strategy which also aligns with the model’s performance reported in Table 4, showing the rationality of our proposed uncertainty-aware self-reflection methods.
5.2 Robustness across Models and Domains
Various Foundation Models
Although Selective Alpaca achieved impressive improvements in LLaMA-2, applying it to other foundation models remains a challenging task. To address this, we apply Selective Alpaca on the Mistral-7B foundation model due to its superior performance Jiang et al. (2023) and present our results on the open-instruct benchmark alignment with the above test configuration. As depicted in Table 4, although Selective Alpaca is selected by the LLaMA-2 models, it is also applicable to the Mistral-7B and improves its capabilities across various tasks, especially on MMLU, BBH, and GSM benchmarks. This experiment fully demonstrates the flexibility of SelectIT which does not rely on a specific foundation model for data selection and the universality of Selective Alpaca which can effectively improve the capabilities of different series or scale LLMs.
Various Domain-specific Tasks
[…]
Table 5.1 shows that SelectIT consistently improves the translation performance of ALMA. Specifically, SelectIT mainly focuses on improving the translation from English to other languages, possibly because ALMA itself has strong English capabilities, which makes it difficult to improve. These results indicate that SelectIT is a versatile method that is effective for IT data selection and applicable to domain-specific tasks.
Method LLaMA-2 ALMA $\Delta$ (normal-$\uparrow$)
7B 13B 7B 13B
Full Dataset 34.3 45.2 29.7 31.5 -
 w/ Random (Full) 34.1 45.1 29.3 31.0 -1.2
 w/ Random (Unselected) 34.6 44.3 29.1 31.2 -1.5
 w/ Length 35.5 47.1 30.1 31.8 +3.8
 w/ SelectIT 36.5 47.7 30.5 32.2 +6.3
Table 6: Comparasion with selection variants.
Figure 4: Instruction embeddings representations of different selection strategies after T-SNE dimensionality reduction. The red and green points are representations of full Alpaca datasets and selected data respectively.
5.3 Insights of Selective Data Curation
Different Selection Strategies
This part compares three different selection strategies, namely, randomly selecting 20% in the full Alpaca, randomly selecting 20% in the unselected dataset, and selecting 20% data based on sample length Zhao et al. (2024), with our SelectIT in the IT and MT experiments. As shown in Table 6, the random-based strategies show certain performance degradation and the random selection in the unselected dataset is even worse, which reflects the effectiveness of our method from the side. Selection based on sample length is a simple approach to defining high-quality data, but it does not take into account the content of IT data, resulting in the limited performance of LLMs. SelectIT can significantly improve the abilities of LLMs.
Data Representation Analysis
This part investigates the relationship between the Selective Alpaca and the original datasets from a representation perspective. We follow Gao et al. (2024) to utilize the outputs of the last layer corresponding to the last token in the input sequence as sample representations. Subsequently, we employ T-SNE (Hinton and Roweis, 2002) for dimensionality reduction, mapping high-dimensional embeddings onto a 2D space. Figure 4 displays the intermediate representations generated by the full and Selective Alpaca respectively. It is observable that it is difficult for randomly selected data to distinguish abnormal data far from the center, making it challenging to define high-quality IT data. However, the majority of Selective Alpaca are concentrated around the central point, indicating that our dataset predominantly contains high-quality data near the center and effectively discards abnormal data distanced from the center which is a finding that strongly supports the conclusion of Table 6.
Data Characteristic Analysis
We analyze the Selective Alpaca from the following two perspectives, to explore why our dataset is better than the original dataset and its variants. Firstly, as shown in Figure  5, the length of instructions from the Selective Alpaca is significantly longer than those in the Alpaca dataset and AlpaGasus which is selected by ChatGPT. This implies that, with the same amount of data, our dataset contains more information, aligned with the results in Table 6. Secondly, by using ChatGPT to examine IT data types, we find a substantial increase in the proportion of computational problems in Selective Alpaca. This indicates that Selective Alpaca tends to select high-quality mathematical data, providing a solid explanation for the observed improvement in the reasoning abilities of LLMs as demonstrated in Table 4. Appendix A.1 shows the case study of comparing the Selective Alpaca with AlpaGasus.
6 Conclusion
Limitation
This paper could be further strengthened as follows:
  • Instruction Data Quantity: Our findings suggest that prioritizing the top 20% of high-quality data optimizes results for Alpaca. Future studies might explore adjusting this threshold based on the data quality in different datasets to enhance performance. Additionally, it would be beneficial to further explore the incorporation of unselected data using curriculum learning (Liu et al., 2020; Zhan et al., 2021; Ding et al., 2021a).
  • Models at Different Scales: Our analysis is currently limited to models smaller than 30B parameters due to computational constraints. Investigating the efficacy of Selective Alpaca on larger-scale LLMs, ranging from 30B to 70B parameters, could provide valuable insights into the method’s scalability and applicability across various model sizes.
  • Diversity of Foundation Models: The study exclusively employs the LLaMA-2 as the foundational model for data selection. Broadening the scope to include alternative foundation models, such as Mistral, may reveal more nuanced data selection strategies and improve overall effectiveness.
  • Expansion to Additional Instruction Datasets: Although SelectIT has been applied to the Alpaca dataset due to its widespread adoption, extending this methodology to incorporate other instruction tuning datasets could offer substantial advantages to the broader LLM research community.
Appendix A Appendix
A.1 Case Study

### v2

@@ 4.2 Main Results
We focus on the discussion of LLaMA-2-13B because both 7B and 13B models exhibit similar trends in Table 4.1. System (10) shows the vanilla IT on LLMs with the original Alpaca. By using the data selection strategies, the ability of LLMs has a moderate enhancement in Systems (12) to (14). Additionally, we can use $S^{base}$ as the input for Equations 4 and 5 to construct individual methods of Sentence-R and Model-R. Systems (15) to (17) illustrate that applying each submodule of SelectIT incrementally enhances LLMs’ performance, rivaling contemporary advanced methods.
Most remarkably, SelectIT can better boost LLaMA-2’s performance compared to vanilla IT in the System (18). Compared to other IT data selection strategies, this enhancement is particularly evident in the computational and reasoning tasks on the BBH and GSM benchmarks. This may be attributed to the characteristics of selected data by SelectIT, and we will analyze this phenomenon in a later section. These gains in reasoning ability also positively impact the coding proficiency of LLMs. The improvement of LLMs on the TydiQA dataset is also obvious enough, which shows that SelectIT can effectively eliminate similar samples and retain sufficient diversity in multilingual aspects.
@@ 5 Analysis
This part aims to answer the research questions through the following experiments: How to select high-quality data in SelectIT? (§5.1) Is SelectIT adaptable to various models and domains? (§5.2) How about the efficiency of SelectIT? (§5.3) What are the advantages of Selective Alpaca?(§5.4)
@@ 5.1 Abalation Study of SelectIT
Figure 3: Comparison of LLM abilities with varying Alpaca proportions.
While SelectIT already excels at assessing and ranking samples effectively, selecting an appropriate number of samples in a redundant dataset remains a crucial aspect of our method. We divide the Alpaca dataset into multiple subsets ranging from 10% to 100% based on SelectIT’s evaluation and evaluate the overall ability of LLMs on the open-instruct benchmark. As illustrated in Figure 3, compared to using the full Alpaca dataset, we observe that LLMs achieve optimal performance using the top 20% to 40% data. Hence, considering the tradeoff of training resources, training time, and model performance, we opt for 20% for implementing the SelectIT on the Alpaca dataset.
@@
3 35.6 46.4 40.5
5 36.2 47.1 41.7
7 35.7 47.3 41.5
9 36.0 46.8 41.4
Table 2: Effect of different $K$.
Effect of Multiple Rating Prompts
$K$ is a critical parameter for our method, impacting not only the range of scores assigned by the LLMs but also the number of rating prompts. We set $K={3,5,7,9}$ and apply SelectIT for sample selection within the Alpaca to get different subset datasets. Table 2 indicates that variations in the value of $K$ have a minor impact on the overall performance of the LLMs. This is attributed to our multi-granularity self-reflection mechanism, which effectively enhances the accuracy and stability of sample selection. Although the model achieves competitive performance at $K$ values of 5 and 7, to minimize resource consumption, we set $K=5$ as the default value in SelectIT.
$\alpha$ MMLU BBH GSM Tydiqa CodeX AE AVG
0.2 47.4 40.6 16.8 47.4 29.4 35.7 36.2
0.4 47.9 39.4 15.5 46.5 29.4 35.8 35.8
0.6 47.8 39.8 16.5 45.6 29.1 35.1 35.7
0.8 47.6 36.4 16.5 43.6 26.7 35.4 34.4
Table 3: Effect of different $\alpha$.
Effect of Uncertainty
$\alpha$ is an uncertainty factor, integral to calibrating the equilibrium between the mean and the standard deviation of scores derived from Token-R. We assign $\alpha$ four different values, i.e., 0.2, 0.4, 0.6, and 0.8, and incorporate SelectIT for sample selection from within Alpaca to generate disparate subset datasets, with all other parameters remaining constant. As shown in Table 3, with a rise in the $\alpha$ value, Sentece-R tends to emphasize the uncertainty innate to LLMs. This results in the neglect of the average score, a fundamental indicator of sample quality, thereby contributing to a decrease in the overall performance of LLMs. Consequently, we ascertain that an $\alpha$ value of 0.2 is optimally suited to establish an effective balance between the sample’s quality and the model’s uncertainty.
ID Individual Unique (%) Overall (%)
@@
Table 4: The relationship between the SelectIT and the individual selection strategy. Sentence-R plays the most significant impact on the final rating of the IT data. IDs 6, 7, and 8 correspond to the system of the same IDs in Table 4.1.
Effect of Different Reflection Strategy
We analyze the relationship between individual selection strategies and SelectIT, from the following two aspects. We first account for the number of high-quality data that can only be selected by a unique selection strategy, referred to as unique selection. Secondly, we calculate which samples in Selective Alpaca can be selected by individual selection strategies in Selective Alpaca, called overall selection. As shown in Table 4, Sentence-R plays the most important role in the final SelectIT strategy. This is because rating prompts play an important role in sample evaluation and exploiting the effect of different prompts on LLM can effectively better improve the accuracy of sample evaluation than Token-R and Model-R. Additionally, this phenomenon also aligns with the model’s performance reported in Table 4.1, showing the rationality of our proposed uncertainty-aware self-reflection methods.
Base Model Datasets Data Size MMLU BBH GSM Tydiqa CodeX AE Overall
AVG $\Delta$ ($\uparrow$)
LLaMA-2-7B LIMA 1K 45.4 37.5 14.3 45.1 24.6 33.1 33.3 -
Selective Alpaca 1K 46.6 41.3 14.5 46.2 30.6 33.8 35.5 +2.2
AlpaGasus 9K 45.9 39.0 14.5 46.4 27.5 35.4 34.8 -
Selective Alpaca 9K 47.2 41.3 18.5 47.6 28.3 35.4 36.4 +1.6
Table 5: Results on IT for different datasets with the same number of instances.
Effect of Data Imbalance
To eliminate unfair comparison caused by IT data quantity imbalance, we adjust the size of the Selective Alpaca dataset to 1,000 and 9,229 respectively, aligning with the LIMA (Zhou et al., 2023) and AlpaGasus (Chen et al., 2024) datasets. The results in Table 5 show that, when facing the same amount of data, SelectIT can still demonstrate better performances, which further illustrates its effectiveness.
5.2 Robustness across Models, Datasets and Domains
Various Foundation Models
Although Selective Alpaca achieved impressive improvements in LLaMA-2, applying it to other foundation models remains a challenging task. To address this, we apply Selective Alpaca on the Mistral-7B and LLaMA-3-8B LLMs and present our results on the open-instruct benchmark alignment with the above test configuration. As depicted in Table 6, although Selective Alpaca is selected by the LLaMA-2 models, it is also applicable to the Mistral-7B, LLaMA-3-8B and improves their capabilities across various tasks, especially on MMLU, BBH, and GSM benchmarks. This experiment fully demonstrates the flexibility of SelectIT which does not rely on a specific foundation model for data selection and the universality of Selective Alpaca which can effectively improve the capabilities of different series or scale LLMs.
Base Model Datasets MMLU BBH GSM Tydiqa CodeX AE Overall
AVG $\Delta$ ($\uparrow$)
LLaMA-2-7B Alpaca-GPT4 46.5 38.4 15.0 43.4 26.8 34.2 34.1 -
Selective Alpaca 47.4 40.6 16.8 47.4 29.4 35.7 36.2 +2.1
LLaMA-2-13B Alpaca-GPT4 55.7 46.6 30.5 47.1 38.8 46.5 44.2 -
Selective Alpaca 55.3 48.5 32.5 54.1 41.2 47.8 46.6 +2.4
@@ 5.2 Robustness across Models, Datasets and Domains
LLaMA-3-8B Alpaca-GPT4 59.6 52.3 34.5 43.1 60.2 48.2 49.7 -
Selective Alpaca 61.2 55.0 37.5 41.1 65.4 47.7 51.3 +1.6
Table 6: Results of IT with various foundation models.
Various Instruction Tuning Datasets
We further validate the robustness of SelectIT by deploying it on two additional, widely-utilized datasets: WizardLM (Xu et al., 2023) and Orca-GPT4 (Subhabrata & Arindam, 2023). WizardLM introduces an innovative method of using LLMs to auto-generate open-domain instructions of varying complexities. This allows for a controlled variation in instructional difficulty and the dataset comprises 143K samples. Orca-GPT4 on the other hand, leverages rich signals from GPT-4 that include explanation traces, step-by-step thought processes, and other multifaceted instructions, all under the guidance of teacher assistance from ChatGPT. Additionally, we maintain consistent hyperparameters, such as $\alpha$ and $K$, choosing LLaMA-2-7B as our base model. We limit the fine-tuning of these datasets to one epoch. As shown in Table 7, SelectIT consistently enhances the performance of the model on both the WizardLM and Orca-GPT4 datasets. Notably, this augmentative effect is especially pronounced in the computational and reasoning tasks within the BBH and GSM benchmarks. In evaluating three separate IT datasets, specifically Alpaca-GPT4, WizardLM, and the more extensive Orca-GPT4, our extensive experimental conclusions validate the broad utility and durability of SelectIT.
Various Domain-specific Tasks
Machine translation (MT) is a representative domain-specific task of LLMs. Previous works have already demonstrated significant improvements with LLMs, but they usually use redundant translation IT datasets. This part tests the robustness of SelectIT on the IT dataset of MT. We select the powerful MT LLM ALMA (Xu et al., 2024) as our backbone model.
Datasets Data Size MMLU BBH GSM Tydiqa CodeX AE Overall
AVG $\Delta$ ($\uparrow$)
WizardLM 143K 43.8 37.8 10.0 41.2 25.2 35.3 32.2 -
WizardLM + SelectIT 28.6K 45.1 40.1 11.0 43.1 27.5 34.7 33.6 +1.4
Orca-GPT4 1M 40.1 35.6 13.0 46.0 23.3 38.1 32.7 -
Orca-GPT4 + SelectIT 0.2M 43.9 38.7 16.5 42.0 27.7 37.4 34.4 +1.7
Table 7: Results of IT with various IT datasets.
[…]
Method Size ALL
COMET BLEU
SoTA Models
NLLB (Costa-jussà et al., 2022) 54B 78.8 26.3
GPT-3.5 - 85.6 34.8
GPT-4 - 85.8 35.1
[…]
LLaMA-2 (Touvron et al., 2023b) 7B 76.5 21.1
TIM (Zeng et al., 2023) 7B 79.1 26.4
SWIE (Chen et al., 2023b) 7B 80.6 27.6
BigTranslate (Yang et al., 2023) 13B 78.8 21.9
Bayling (Zhang et al., 2023) 13B 82.0 27.8
[…]
ALMA 7B 83.2 29.7
   w/ SelectIT 7B 83.7 30.5
[…]
   w/ SelectIT 13B 84.2 32.2
Table 8: The overall results on MT LLMs.
examples were selected using SelectIT. We utilize both BLEU (Post, 2018; Ott et al., 2018) and COMET (Rei et al., 2022) based on the wmt22-comet-da model for evaluation. We report results for the two language pairs in four directions, using ALL to represent their average. Table 5.2 shows that SelectIT consistently improves ALMA’s translation performance. These results indicate that SelectIT is a versatile and scalable method, effective not only for IT data selection but also for domain-specific tasks like MT. For more detailed analysis and results, please see Appendix A.1.
5.3 Efficiency of SelectIT
Method Speed Time Cost
ChatGPT API 0.76 it/s 19.07h \$52.02
GPT4 API 0.37 it/s 38.98h \$2871.56
SelectIT 9.34 it/s 5.80h \$26.68
Table 9: Comparison of selection efficiency.
SelectIT is a faster and more cost-effective method for IT data selection. We compared different selection methods on the Alpaca-GPT4 dataset. For ChatGPT (AlpaGasus) or GPT-4, we randomly select 500 instruction data from Alpaca-GPT4, analyze various metrics, and estimate the resource consumption for selecting the entire dataset. Using SelectIT, we employ 4 A800 80G GPUs to select high-quality IT data, calculating the total cost based on Google Cloud’s rate of \$1.15/h per single GPU. As shown in Table 9, SelectIT is significantly faster and uses the least resources. This efficiency is due to computing only the probability of the next token for input sentences, bypassing the full sentence generation and decoding process, resulting in lower resource consumption. Additionally, using our own GPU at a low cost enhances transparency, allowing us to preserve all intermediate outputs and results for thorough analysis in data selection.
5.4 Insights of Selective Data Curation
Different Selection Strategies
Method LLaMA-2 ALMA $\Delta$ ($\uparrow$)
7B 13B 7B 13B
Full Dataset 34.1 44.2 29.7 31.5 -
   w/ Random (Full) 34.1 45.1 29.3 31.0 0.0
   w/ Random (Unselected) 34.6 44.3 29.1 31.2 -0.4
   w/ Length 35.5 47.1 30.1 31.8 +5.0
   w/ SelectIT 36.2 47.1 30.5 32.2 +6.5
Table 10: Comparasion with variants.
This part compares three different selection strategies, namely, randomly selecting 20% in the full Alpaca and unselected dataset of Selective Alpaca, and selecting 20% data based on sample length (Zhao et al., 2024). As shown in Table 10, the random-based strategies show certain performance degradation and the random selection in the unselected dataset is even worse, which reflects the effectiveness of our method from the side. Selection based on sample length is a simple approach to defining high-quality data, but it does not take into account the content of IT data, resulting in the limited performance of LLMs. SelectIT can significantly improve the abilities of LLMs.
Data Representation Analysis
This part explores the relationship between Selective Alpaca and the original datasets from a representation perspective. Following Gao et al. (2024), we use the outputs of the last layer corresponding to the last token in the input sequence as sample representations. We then apply T-SNE (Hinton & Roweis, 2002) for dimensionality reduction, mapping high-dimensional embeddings onto a 2D space. Figure 5 shows the intermediate representations generated by the full and Selective Alpaca datasets. Randomly selected data struggle to distinguish abnormal data far from the center, making it hard to define high-quality IT data. In contrast, Selective Alpaca data are mostly concentrated around the center, indicating that our dataset predominantly contains high-quality data near the center and effectively discards abnormal data, supporting the conclusion of Table 10.
Figure 4: Instruction embeddings representations of different selection strategies. The red and blue points are representations of full Alpaca datasets and selected data respectively.
Data Characteristic Analysis
We analyze the Selective Alpaca from the following two perspectives, to explore why our dataset is better than the original dataset and its variants. Firstly, as shown in Figure 5, the length of instructions from the Selective Alpaca is significantly longer than those in the Alpaca dataset and AlpaGasus which is selected by ChatGPT. This implies that, with the same amount of data, our dataset contains more information, aligned with the results in Table 10. Secondly, by using ChatGPT to examine IT data types, we find a substantial increase in the proportion of computational problems in Selective Alpaca. This indicates that Selective Alpaca tends to select high-quality mathematical data, providing a solid explanation for the observed improvement in the reasoning abilities of LLMs as demonstrated in Table 4.1. Appendix A.3 shows the case study of comparing the Selective Alpaca with AlpaGasus.
Figure 6: Changing trends of the calculation and sample length with different data sizes.
Insights of High-Quality Data in SelectIT
Furthermore, we analyze the proportion of calculation and sample average length in Alpaca-GPT4 with different proportions after sorting by SelectIT to explore its intrinsic characteristics and the definition of high-quality data. As shown in Figure 6, with the proportion of Alpaca-GPT4 data continuing to increase, the proportion of calculation and sample average length gradually decreases. This phenomenon clearly indicates that SelectIT can reasonably rank samples based on their characteristics. When the data size is more than 50%, the proportion of calculation IT data sharply declines, falling below 6%, causing a noticeable decrease in the model’s overall capability, as depicted in Figure 3. This analysis shows that more computationally intensive IT data may be a new perspective on the characteristics of optimal IT data, which not only effectively improves the LLMs’ reasoning ability, but also further drives the improvement of other abilities.
6 Conclusion
Limitation
This paper could be further strengthened as follows:
  • Instruction Data Quantity: Our findings suggest that prioritizing the top 20% of high-quality data optimizes results for Alpaca. Future studies might explore adjusting this threshold based on the data quality in different datasets to enhance performance.
  • Models at Different Scales: Our analysis is currently limited to models smaller than 30B parameters due to computational constraints. Investigating the efficacy of Selective Alpaca on larger-scale LLMs, could provide valuable insights into the method’s scalability.
  • Expansion to Additional Instruction Datasets: Although SelectIT has been applied to the Alpaca dataset due to its widespread adoption, extending this methodology to incorporate other IT datasets could offer substantial advantages to the broader LLM research community.
Broader Impacts
Our work follows the NeurIPS Ethics Policy. Our findings are based on publicly available datasets for reproducibility purposes. LLMs can contain potential racial and gender bias. Therefore, if someone finds our work interesting and would like to use it in a specific environment, we strongly suggest the user check the potential bias before usage. In addition, it is hard to control the generation of LLMs. We should be aware of the potential problems caused by hallucinations.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62206076), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515011491), Shenzhen Science and Technology Program (Grant Nos. ZDSYS20230626091203008, KJZD20231023094700001, RCBS20221008093121053), and Shenzhen College Stability Support Plan (Grant Nos. GXWD20220811173340003, GXWD20220817123150002). Derek F. Wong was supported in part by the Science and Technology Development Fund, Macau SAR (Grant Nos. FDCT/060/2022/AFJ, FDCT/0070/2022/AMJ), National Natural Science Foundation of China (Grant No. 62261160648), the Multi-year Research Grant from the University of Macau (Grant No. MYRG-GRG2024-00165-FST), and the Tencent AI Lab Rhino-Bird Gift Fund (Grant No. EF2023-00151-FST). We would like to thank the anonymous reviewers and meta-reviewer for their insightful suggestions.
Appendix A Appendix
A.1 Applying SelectIT on Machine Translation LLMs
Machine Translation (MT) is a important task for LLMs, demonstrating their domain-specific capabilities. Prior research, including TIM (Zeng et al., 2023), SWIE (Chen et al., 2023b), BigTranslate (Yang et al., 2023), and Bayling (Zhang et al., 2023), has shown significant improvements in LLMs, often relying on extensive translation training datasets. In this section, we examine the impact of training data quality on MT performance, employing the robust MT LLM, ALMA, as our foundational model (Xu et al., 2024).
[…]
We report both the widely used BLEU score (Post, 2018; Ott et al., 2018) and the COMET score (Rei et al., 2022) based on the wmt22-comet-da model, which shows higher correlation with human judgments for evaluating the LLMs’ translation abilities. Table A.1 consistently demonstrates that SelectIT enhances ALMA’s translation efficacy. Notably, SelectIT primarily focuses on improving translations from English to other languages, likely due to ALMA’s inherent proficiency in English, which presents challenges for further enhancements. These findings highlight SelectIT’s adaptability and scalability, validating its effectiveness not only in IT data selection but also in domain-specific tasks such as MT.
@@ A.1 Applying SelectIT on Machine Translation LLMs
Table 11: Overall results on machine translation LLMs. “” the improvement is significant by contrast to the ALMA model ($p$ < 0.05).
Figure 7: Example on Sentence-R calculation of SelectIT.
A.2 Details of Sentence-level Rating
Based on the preceding analysis, Sentence-R is integral to the functionality of SelectIT. As illustrated in Equation 4, the Token-level Rating forms the foundation for the Sentence-level Rating. The Model-level Rating is derived through multiple iterations of the Sentence-level Rating across different foundational LLMs. Therefore, a detailed explanation of Sentence-R is sufficient to demonstrate the operational mechanism of SelectIT. As depicted in Figure 7, we utilize five distinct rating prompts along with a single input to formulate the final input for Sentence-R. Initially, each rating prompt produces a score of $S^{token}$. We then compute the mean and standard deviation of these $S^{token}$ values to obtain the final $S^{sent}$, as outlined in Equation 4.
A.3 Case Study